OpenStudy (anonymous):

what is the wavelength in nm of radiation that has an energy content of 1.0x10to the power of 3 kj/mol? in which region of the electromagnetic spectrum is this radiation found in? (can u please explain step-by-step) thank you

OpenStudy (wolfe8):

Hmm is that all the info you have?

OpenStudy (wolfe8):

No frequency?

OpenStudy (aaronq):

don't need the frequency. \(E_{photon}=\dfrac{h*c}{\lambda} \rightarrow \lambda=\dfrac{h*c}{E_{photon}}\) h = plancks constant, c= speed of light

OpenStudy (wolfe8):

Ah so resolve the frequency to c/wavelength then. Good call!